OpenStudy (anonymous):

Does the series of cos (1/n) converge or diverge ? please explain it would be really helpful.

OpenStudy (anonymous):

cosine is continuous right?

OpenStudy (anonymous):

i assume you mean "sequence" here, not series unless you are adding

OpenStudy (anonymous):

so \[\lim_{n\to \infty}\cos(\frac{1}{n})=\cos(\lim_{n\to \infty}\frac{1}{n})=\cos(0)=1\]

OpenStudy (anonymous):

if it is in fact \[\sum_{n=1}^{\infty}\cos(\frac{1}{n})\] then that is a totally different story, and i would have to do some thinking

OpenStudy (anonymous):

oh, no i wouldn't the terms do not go to zero, they go to one so no chance it converges

OpenStudy (anonymous):

it is the sum of of cos (1/n)

OpenStudy (anonymous):

ok then see answer above the terms do not even go to zero, they get closer and closer to 1 so the sum cannot converge

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

Yea I understand. Thanks so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i have no idea?.? sorrry